1.27 μm O_{2} CONTINUUM ABSORPTION IN O_{2}/CO_{2} MIXTURES

G.T. FRASER and W.J. LAFFERTY, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441, USA.

The collision-induced, near-infrared O_{2} continuum band overlapping the weak a^{1}Δ_{g} − X^{3}Σ_{g}^{−} , v = 0 - 0, 1.27 μm discrete band of O_{2} has been investigated in O_{2}/CO_{2} mixtures at room temperature (T = 296 K) for total densities from 1.8 to 9.3 times that of an ideal gas under standard conditions (T = 273.15 K and P = 101.325 kPa), i.e., from 1.8 to 9.3 amagats. Absorption spectra were recorded at 0.5 cm^{−1} resolution using a Fourier-transform spectrometer and an 84-m pathlength. A least-squares analysis of the integrated band strength, S_{total} = S_{O2}ρ_{O2} + S_{O2−O2}ρ_{O2}^{2} + S_{O2−CO2}ρ_{O2}ρ_{CO2}, as a function of the carbon dioxide density, ρ_{CO_{2}}, and the oxygen density, ρ_{O_{2}}, yields S_{O_{2}−CO_{2}} = 2.95(40) × 10^{−43} cm^{−2}(molecule/cm^{3})^{−2} [i.e., 2.13(29) × 10^{−4} cm^{−2} amagat^{−2}]. The S_{O2−CO2} coefficient is approximately three times greater than the corresponding S_{O2−N2} coefficient determined from studies of O_{2}/N_{2} mixtures, illustrating the efficiency of large electric multipolar moments in inducing continuum absorption in the 1.27 μm band of O_{2}. A similar large enhancement of the O_{2} continuum absorption by CO_{2} is observed for the v = 1 - 0, O_{2} vibrational fundamental. The results support the calculations by Brown and Tipping, which demonstrate the importance of water, with its large electric dipole moment, in enhancing the collision-induced absorption bands of O_{2} and N_{2} in the atmosphere. We suggest that the apparent inability of radiative-transfer models to accurately account for the increased atmospheric absorption present when water-vapor levels increase may be due in part to the neglect of the intensity enhancement of a number of continuum bands and of the far wings of discrete bands by water-vapor collisions.